OpenStudy (anonymous):

For the given quadratic equation convert into vertex form, find the vertex, and find the value for x = 6. y= -2x^2 + 2x +2?

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

if we make this substitution: x=6 into your quadratic equation, we get: \[\Large \begin{gathered} y = - 2 \times {6^2} + 2 \times 6 + 2 = \hfill \\ \hfill \\ = - 2 \times 36 + 12 + 2 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

-58

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

So that's the value of x, correct?

OpenStudy (michele_laino):

that's the value of y, when x=6

OpenStudy (anonymous):

So how do we find the value of x and find the vertex?

OpenStudy (michele_laino):

our task was to find the value of y, when x=6

OpenStudy (michele_laino):

the general formula of a parabola is: \[\Large y = a{x^2} + bx + c\]

OpenStudy (anonymous):

So there is nothing else to the problem it's done that simple lol

OpenStudy (michele_laino):

we have to find the coordinates of the vertex, and the vertex form of the equation of your parabola

OpenStudy (michele_laino):

so, as I wrote before, the general equation of a parabola, whose axis is parallel to y axis, is: \[\Large y = a{x^2} + bx + c\] now, please by comparison with the equationof your parabola, what are the coefficients a, b, and c?

OpenStudy (anonymous):

6 and -58?

OpenStudy (michele_laino):

hint: